When 50.0 g of 0.200 M NaCl(aq) at 24.1 °C is added to 100.0 g of 0.100 M AgNO3(aq) at 24.1 °C in a
calorimeter, the temperature increases to 25.2 °C as AgCl(s) forms. Assuming the specific heat of the solution and
products is 4.20 J/g °C, calculate the approximate amount of heat in joules produced.
Answer #1

Q = m*C*delta T

m = mass of the soluton

C = specific heat

delta T = Temperature change

Q = heat released

Q = 150*4.2*1.1 = 693 Joules

693 joules of heat is produces